OpenStudy (anonymous):

(4c^5+11c^2-4)-(8c^5-9c^2+18)

OpenStudy (compassionate):

\[(4c^5+11c^2-4)-(8c^5-9c^2+18) \rightarrow 4c^5 - 11c^2 - 4 - 8c^5 + 9c^2 - 18\] What I did above was multiply everything in the second parenthesis by negative -. Remember, when a negative sign is on the outside, all the signs within the parenthesis change. Now, you need to add like terms. I'll group them together to make it easier. \[(4c^5 - 8c^5) (- 11c^2+ 9c^2) (-4 - 18)\] NOTE: we wouldn't really need to put the parenthesis. This is just for visualization so you can see the like-terms. Now, just add them like normal. Can you do the final step?

OpenStudy (anonymous):

-4c^5-2c^2-12

OpenStudy (anonymous):

Am i correct

OpenStudy (compassionate):

Close! \[-4c^5 - 2c^2 - 22\] What you did was get -12. It would only be -12 if it was: -4 -(-14) A double minus = positive But just having: - 4 - 18, you still get a negative. Also: \(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) Study on ~~

OpenStudy (anonymous):

Thank you, do you mind if you help me with more similar questions

OpenStudy (compassionate):

No problem. Remember, if you want to ask a new question, open a Ask a Question. Don't ask two questions on the same thread! :) And always give a medal to someone who you found helpful.